# Approximating Robot Configuration Spaces with few Convex Sets using Clique Covers of Visibility Graphs

### v1

@@ I Introduction
Approximating complex sets as a union of simpler sets is a common pre-processing method for accelerating downstream computations. Familiar examples include clustering methods for high-dimensional data in machine learning [1, §2.5.2], approximating complex shapes via triangular meshes to facilitate graphics rendering [2], and describing geometries as unions of convex sets for efficient collision checking [3]. Similarly, recently developed methods for robot motion planning rely on (conservative) decompositions of the environment into convex sets to design smooth trajectories around obstacles using efficient convex optimization [4, 5, 6, 7, 8]. These motion planners have demonstrated great potential, and performance frequently superior to widely used sampling-based methods [5]. However, the decomposition of the environment that they require is a daunting task, often demanding a substantial degree of human supervision. This is largely due to the fact that the collision-free subset of a robot’s configuration space ($\mathcal{C}^{\mathrm{free}}$) is intractable to describe analytically [9, §3], even if the robot’s task space is relatively simple (see Figure 1 for an example).
@@
Figure 1: The collision-free configuration space of a simple robot is decomposed into 7 polytopes, achieving around 92% coverage. Left: Robot with with 3 revolute joints $q_{1}$ to $q_{3}$. Center: Visualization of the full collision-free configuration space $\mathcal{C}^{\mathrm{free}}$, given by the interior of the green mesh. Right: Approximate convex cover of $\mathcal{C}^{\mathrm{free}}$ generated with the proposed method. See also: Interactive visualization, video.
@@
Through a large variety of experiments, we show that our algorithm outperforms the approaches that have been previously used in the literature to seed and inflate convex regions; both in terms of runtimes and number of regions used to cover the space. As an example, for a robot arm with seven degrees of freedom, and many task-space obstacles, our method requires approximately $46$ regions and one hour of computations to cover $70\%$ of $\mathcal{C}^{\mathrm{free}}$. Whereas the approach outlined in [12] requires ten times more regions and is ten times slower.
@@ II-A Algorithms Requiring Explicit Obstacle Descriptions
The recent work [18] constructs low-cardinality convex covers of two-dimensional polygons by first decomposing them into small convex pieces (for example triangles), and then by joining the pieces based on a small clique cover of an approximated set-to-set visibility graph. In three dimensions the approach in [19] can be used to decompose $\mathcal{C}^{\mathrm{free}}$ into sets that are approximately convex, provided that a triangle mesh description is available. Finally, if the configuration-space obstacles are explicitly described as convex sets, the original IRIS algorithm from [10] can be used to inflate a large polytope in $\mathcal{C}^{\mathrm{free}}$ around a specified seed point in arbitrary dimensions.
@@ II-B Algorithms Allowing Implicit Obstacle Descriptions
Most commonly, the explicit descriptions of the obstacles are only available in task space; while the collision-free configuration space $\mathcal{C}^{\mathrm{free}}$ is defined implicitly through the robot’s inverse kinematics, and intractable to describe analytically [9, §3]. Obtaining convex decompositions in this setting has been the focus of multiple works. The method in [20] uses visibility graphs and kernels to compute convex decompositions of three-dimensional spaces via sample-based collision checking. However, that method represents the convex sets through their vertices, and is inefficient in higher dimensions. In the family of IRIS algorithms, two methods can deal with implicit descriptions of $\mathcal{C}^{\mathrm{free}}$: IRIS-NP [12] and C-IRIS [11, 13]. The former extends the original IRIS method [10] to arbitrary configuration spaces using nonlinear programming and inflates polytopes that are collision-free with high probability. The latter grows polytopes that are rigorously certified to be collision-free using a rational reparametrization of the configuration space and sums-of-squares programming.
@@ III-A Problem Statement
In practice, we are interested in solving this problem for values of $\alpha$ that are sufficiently high to accomplish a task of interest, such as collision-free motion planning, but not so large that the cardinality $N$ grows unreasonably. Indeed, when $\alpha=1$, Min $\alpha$-ApproxConvexCover might not even have a finite solution. [^2] For $\alpha<1$ a finite solution is guaranteed to exist. This can be seen by approximating the volume of $\mathcal{C}^{\mathrm{free}}$ over finite hyperrectangle partitions with a lower Darboux integral.
@@ IV-A Sampling the Visibility Graph
At the beginning of every iteration of VCC, the subroutine SampleVisibilityGraph samples $K$ configurations uniformly at random from the portion of $\mathcal{C}^{\mathrm{free}}$ that is not already covered by the polytopes in $\mathcal{R}$. Then, it constructs the visibility graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, by checking for collisions along the line segments connecting each pair of sampled configurations. Currently, this is performed using sampling-based collision checkers. Exact visibility checking is possible using methods such as [23][24, §5.3.4], or [13].
@@ IV-C Summarizing Cliques with Ellipsoids
In the subroutine MinVolumeEllipsoids, we solve a semidefinite program to enclose each clique with an ellipsoid of minimum volume [25, §8.4.1]. This collection $\mathcal{B}$ of ellipsoids allows us to summarize the geometry of each clique with a point and a set of principal directions, which are then used to initialize the region-inflation algorithm. For the upcoming computations, it is necessary that the center of each ellipsoid is not in collision; if this is not the case, we recenter the ellipsoid around the vertex in the clique that is closest to its center.
@@ IV-D Inflating Polytopes
In the last step of VCC, the subroutine InflatePolytopes inflates a large collision-free polytope around the center of each ellipsoid induced by a clique.
Let us initially assume that the obstacles are convex. Consider a single ellipsoid. Using convex optimization we compute the point in each obstacle that is closest to the center of the ellipsoid, according to the distance metric induced by the ellipsoid. These points anchor separating hyperplanes between the ellipsoid center and the obstacles, which form a polytope of obstacle-free space. These steps are repeated for each ellipsoid (i.e., for each clique) to obtain a collection of collision-free polytopes that we add to the set $\mathcal{R}$.
Domain Village 3DOF Flipper 5DOF UR5 7DOF IIWA
@@ V Experiments
We evaluate VCC and IOS on four environments: Village, 3DOF Flipper, 5DOF UR5, and 7DOF IIWA. The dimension $n$ of these environments ranges from $3$ to $7$. For each environment, we run the two algorithms ten times and report their performance in Table I. The Village environment from [8] contains only convex obstacles and is compatible with the original IRIS [10]. All other examples involve the configuration spaces of robotic manipulators and therefore IRIS-NP [12] is employed. The number of samples used in the convergence check in Algorithm 1 is $M=5000$.
VCC meets the required coverage threshold with significantly fewer regions and in a substantially shorter amount of time. Notably, in the most challenging benchmark, 7DOF IIWA, VCC requires 10 times fewer regions and meets the required coverage of $70\%$ around 16 times faster. This substantial speedup can be attributed to two key factors. First, the region inflation in VCC is parallelized. Second, IRIS is the most computationally expensive step. While VCC only requires a single iteration of IRIS per region, IOS can require up to around five IRIS iterations per region, due to the initialization with a potentially uninformative spherical metric.
@@ VI Limitations of Approximating Convex Sets with Cliques
$c_{i}^{T}q_{j}+d_{i}$ $\leq M(1-b_{j}),\quad$ $\forall i,j=1,\dots,K,$ (2b)
where $M$ is a large enough constant (e.g. the maximum distance between all pairs of vertices).

### v2

@@ I Introduction
Approximating complex sets as a union of simpler sets is a common pre-processing method for accelerating downstream computations. Familiar examples include clustering methods for high-dimensional data in machine learning [1, §2.5.2], approximating complex shapes via triangular meshes to facilitate graphics rendering [2], and describing geometries as unions of convex sets for efficient collision checking [3]. Similarly, recently developed methods for robot motion planning rely on (conservative) decompositions of the environment into convex sets to design smooth trajectories around obstacles using efficient convex optimization [4, 5, 6, 7, 8]. These motion planners have demonstrated great potential, and performance frequently superior to widely used sampling-based methods [5]. However, the required decomposition of the environment is a daunting task, often demanding a substantial degree of human supervision. This is largely due to the fact that the collision-free subset of a robot’s configuration space ($\mathcal{C}^{\mathrm{free}}$) is intractable to describe analytically [9, §3], even if the robot’s task space is relatively simple (e.g. see Figure 1).
@@
Figure 1: The collision-free configuration space of a simple robot is decomposed into 7 polytopes, achieving around 92% coverage. Left: Robot with with 3 revolute joints $q_{1}$ to $q_{3}$. Center: Visualization of the full collision-free configuration space $\mathcal{C}^{\mathrm{free}}$, given by the interior of the green mesh. Right: Approximate convex cover of $\mathcal{C}^{\mathrm{free}}$ generated with the proposed method. See also: Website, video.
@@
Through a large variety of experiments, we show that our algorithm outperforms previous approaches for seeding and inflating convex regions; both in terms of runtimes and number of regions used to cover the space. As an example, for a robot arm with seven degrees of freedom, and many task-space obstacles, our method requires approximately $46$ regions and one hour of computations to cover $70\%$ of $\mathcal{C}^{\mathrm{free}}$, whereas the approach outlined in [12] requires ten times more regions and is ten times slower.
@@ II-A Algorithms Requiring Explicit Obstacle Descriptions
The recent work [18] constructs low-cardinality convex covers of two-dimensional polygons by first decomposing them into small convex pieces (for example triangles), and then by joining the pieces based on a small clique cover of an approximated set-to-set visibility graph. Similarly, the visibility graph of polygons is directly used to construct convex covers in [19]. In three dimensions the approach in [20] can be used to decompose $\mathcal{C}^{\mathrm{free}}$ into sets that are approximately convex, provided that a triangle mesh description is available. Finally, if the configuration-space obstacles are explicitly described as convex sets, the original IRIS algorithm from [10] can be used to inflate a large polytope in $\mathcal{C}^{\mathrm{free}}$ around a specified seed point in arbitrary dimensions.
@@ II-B Algorithms Allowing Implicit Obstacle Descriptions
Most commonly, the explicit descriptions of the obstacles are only available in task space; while the collision-free configuration space $\mathcal{C}^{\mathrm{free}}$ is defined implicitly through the robot’s inverse kinematics, and is intractable to describe analytically [9, §3]. Some works have focused on approximately describing the configuration space obstacles, rather than $\mathcal{C}^{\mathrm{free}}$ [21, 22, 23]. Conversely, multiple works have studied directly obtaining convex decompositions of $\mathcal{C}^{\mathrm{free}}$. In [24] visibility graphs and kernels are used to compute convex decompositions of three-dimensional spaces via sample-based collision checking. However, their method represents the convex sets through their vertices, and is inefficient in higher dimensions. In the family of IRIS algorithms, two methods can deal with implicit descriptions of $\mathcal{C}^{\mathrm{free}}$: IRIS-NP [12] and C-IRIS [11, 13]. The former extends the original IRIS method [10] to arbitrary configuration spaces using nonlinear programming and inflates polytopes that are collision-free with high probability. The latter grows polytopes that are rigorously certified to be collision-free using a rational reparametrization of the configuration space and sums-of-squares programming.
@@ III-A Problem Statement
In practice, we are interested in solving this problem for values of $\alpha$ that are sufficiently high to accomplish a task of interest, such as collision-free motion planning, but not so large that the cardinality $N$ grows unreasonably. Indeed, when $\alpha=1$, Min $\alpha$-ApproxConvexCover might not even have a finite solution. [^2] For $\alpha<1$ a finite solution is guaranteed to exist. This can be seen by computing the volume of $\mathcal{C}^{\mathrm{free}}$ with a lower Darboux integral over finite hyperrectangle partitions.
@@ IV-A Sampling the Visibility Graph
At the beginning of every iteration of VCC, the subroutine SampleVisibilityGraph samples $K$ configurations uniformly at random from the portion of $\mathcal{C}^{\mathrm{free}}$ that is not already covered by the polytopes in $\mathcal{R}$. Then, it constructs the visibility graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, by checking for collisions along the line segments connecting each pair of sampled configurations. Currently, this is performed using sampling-based collision checkers. Exact visibility checking is possible using methods such as [27][28, §5.3.4], or [29].
@@ IV-C Summarizing Cliques with Ellipsoids
In the subroutine MinVolumeEllipsoids, we solve a semidefinite program to enclose each clique with an ellipsoid of minimum volume [30, §8.4.1]. This collection $\mathcal{B}$ of ellipsoids allows us to summarize the geometry of each clique with a point and a set of principal directions, which are then used to initialize the region-inflation algorithm. For the upcoming computations, it is necessary that the center of each ellipsoid is not in collision; if this is not the case, we recenter the ellipsoid around the vertex in the clique that is closest to its center.
@@ IV-D Inflating Polytopes
In the last step of VCC, the subroutine InflatePolytopes inflates a large, collision-free polytope around the center of each ellipsoid induced by a clique.
Let us initially assume that the obstacles are convex. Consider a single ellipsoid; using convex optimization, we compute the point in each obstacle that is closest to the center of the ellipsoid, according to the distance metric induced by the ellipsoid. These points anchor separating hyperplanes between the ellipsoid center and the obstacles, which form a polytope of obstacle-free space. These steps are repeated for each ellipsoid (i.e., for each clique) to obtain a collection of collision-free polytopes that we add to the set $\mathcal{R}$.
Domain Village 3DOF Flipper 5DOF UR3 7DOF IIWA
@@ V Experiments
We evaluate VCC and IOS on four environments: Village, 3DOF Flipper, 5DOF UR3, and 7DOF IIWA. The dimension $n$ of these environments ranges from $3$ to $7$. For each environment, we run the two algorithms ten times and report their performance in Table I. The Village environment from [8] contains only convex obstacles and is compatible with the original IRIS [10]. All other examples involve the configuration spaces of robotic manipulators and therefore IRIS-NP [12] is employed. The number of samples used in the convergence check in Algorithm 1 is $M=5000$.
VCC meets the required coverage threshold with significantly fewer regions and in a substantially shorter amount of time. Notably, in the most challenging benchmark, 7DOF IIWA, VCC requires 10 times fewer regions and meets the required coverage of $70\%$ around 16 times faster. This substantial speedup can be attributed to two key factors. First, the region inflation in VCC is parallelized. Second, IRIS is the most computationally expensive step. While VCC only requires a single iteration of IRIS per region, IOS can requires around five IRIS iterations per region, due to the initialization with a potentially uninformative spherical metric.
@@ VI Limitations of Approximating Convex Sets with Cliques
$c_{i}^{T}q_{j}+d_{i}$ $\leq L(1-b_{j}),\quad$ $\forall i,j=1,\dots,K,$ (2b)
where $L$ is a large enough constant (e.g. a constant factor times the diameter of the visibility graph).